OpenStudy (anonymous):

A 20ft ladder is leaning against a building. If the base of the ladder is 6ft from the base of the building, what is the angle of elevation of the ladder? How high the ladder reach on the building?

OpenStudy (zzr0ck3r):

do you know soh cah toa?

OpenStudy (anonymous):

Yes i've heard of it, but i don't really know how to use it. :/

OpenStudy (zzr0ck3r):

what are you doing in class?

OpenStudy (anonymous):

Well we just started learning about Sin, Cos, Tan, and how to use them with right triangles.

OpenStudy (zzr0ck3r):

soh says sin(x) = opposite over hypotenuse cah says cos(x) = adjacent over hypotenuse toa says tan(x) = opposite over adjacent

OpenStudy (zzr0ck3r):



OpenStudy (zzr0ck3r):

so you are given the hypotonuse (20) and adjacent (6) so cos(x) = 20/6 = 10/3

OpenStudy (anonymous):

Wow, that was actually pretty simple. Thank you very much :)

OpenStudy (anonymous):

But why do we only have to find the cos(x) and not the other ones?

OpenStudy (zzr0ck3r):

well it depends on what you are given. If they told you adjacent and opposite then you would use toa so tan(x) = opposite/adjacent

OpenStudy (zzr0ck3r):

so if they told you the hight and 6ft bit, you would use tan

OpenStudy (anonymous):

Thank you (:

OpenStudy (zzr0ck3r):

you still need the height, use a^2+b^2 = c^2 you are given a and c so solve for b

OpenStudy (anonymous):

oh i remember how to do that !! (: Thanks again.